Calculate pI value of Asp (pka of COOH is 1.99 ,pka of NH3+ is 9.90 ,pKa of R group is 3.90)
Answer #1

pI value of Asp ( aspartic acid ) is 2.945.

We have drawn structure of aspartic acid now the pI of aspartic acid of average of pKa1 and pKa2. Which is 2.945.

Note:-

If simple amino acid with non ionizable side chain so pI is the average of given two pKa values.

If R group is acidic or if it is acidic amino acid then the pI is the average of two lowest pKa values.

It R group is basic or basic amino acid present then pI is the average of two highest pka values.

,K Pkaj = 1.99 - R CH₂ Cool Cook < Pkag= 3.90 H₂N-GH-Cool CH₂ group Pkaz= 9.90 (Aspartic acid) gf dide group R is acidic so I